# Graph Sampling for Scalable and Expressive Graph Neural Networks on Homophilic Graphs

### v3

@@ IV Experimental Results
In this section, we present an empirical analysis of our sampling algorithm on citation networks. Specifically, we compare the test accuracy achieved by GNN models trained on subgraphs sampled using our heuristic and the random node sampling baseline on the full size graph.
Trace preservation. First, we investigate the ability of our heuristic to preserve Laplacian trace in three homophilic citation networks. In agreement with the lower bound derived in Proposition III.2, in Figure 1 we see that across all datasets, our sampling method results in subgraphs with larger adjusted $\tilde{{tr}}(\textbf{L})={tr}(\mathbf{L})/n$. This pattern is especially visible for PubMed, as exemplified by the 800-node subsamples in Figure 2. Clearly, the graph sampled using Algorithm 1 is more connected than its random counterpart.
GNN training. Next, we investigate the transferability of GNNs trained on subgraphs sampled using our heuristic and subgraphs and random node-induced subgraphs. Specifically, we train a GNN model on graphs of size given by the $x$-axis of Figure 3, and report their test accuracy on the full size graph.
@@
In Figure 3, the box plots represent the mean and standard error of the accuracy achieved over 50 rounds, and the red dots the test accuracy achieved by our model. Our sampling heuristic is substantially better, except for CiteSeer at sample rate 25%. Notably, at a sample rate of 62.5%, our heuristic beat random by around 10% for all three datasets.

### v5

@@ IV Experimental Results
In this section, we present an empirical analysis of our sampling algorithm on the Planetoid citation networks Cora, CiteSeer, and PubMed, with feature homophilies given by -0.0415, -0.0194 and -0.0007 respectively. Specifically, we compare the test accuracy achieved by GNN models trained on subgraphs sampled using our heuristic and the random node sampling baseline on the full size graph.
Trace preservation. First, we investigate the ability of our heuristic to preserve Laplacian trace in three homophilic citation networks. In agreement with the lower bound derived in Proposition III.2, in Figure 1 we see that across all datasets but CiteSeer, our sampling method generally results in subgraphs with larger adjusted $\tilde{{tr}}(\textbf{L})={tr}(\mathbf{L})/n$ on average. This pattern is especially visible for PubMed (the most feature homophilic network), as exemplified by the 800-node subsamples in Figure 2. Clearly, the graph sampled using Algorithm 1 is more connected than its random counterpart.
GNN training. Next, we investigate the transferability of GNNs trained on subgraphs sampled using our heuristic and random node-induced subgraphs. Specifically, we train a GNN model on graphs of size given by the $x$-axis of Figure 3, and report their test accuracy on the full size graph.
Experiment details. For each dataset and sample rate, we chose hidden dimension 128, 2 layers, 300 epochs, learning rate and weight decay 0.001 and 0.0001 respectively, and a GCN architecture with ReLU activations. All were optimized using Adam [kingma17-adam]. All of the graphs are symmetrized beforehand, turning directed edges into undirected ones.
In Figure 3, the error bars show the mean and standard error of the accuracy achieved over 50 rounds, and the red dots the test accuracy achieved by our model. Our sampling heuristic is generally better, especially for PubMed (the most feature homophilic network); and especially at higher sample sizes (fewer deletions) and lower sample sizes (substantial deletions).
Proof of Prop. II.2.
From the definition, we can rewrite
${\mathcal{Y}}=\text{span}(\{{\mathbf{S}}^{k}x:k=0,\ldots,K\})$
@@
Proof of Prop. III.2.
By the definition of feature homophily, and using the Cauchy-Schwarz inequality, we write
$nh_{G}=tr({\mathbf{L}}\hat{X}\hat{X}^{T})=\langle{\mathbf{L}},\hat{X}\hat{X}^{T}\rangle_{F}\leq\|{\mathbf{L}}\|_{F}\cdot\|\hat{X}\hat{X}^{T}\|_{F}\text{.}$ (6)
This means, by squaring both sides,
$n^{2}h_{G}^{2}\leq\|{\mathbf{L}}\|_{F}^{2}\cdot\|\hat{X}\hat{X}^{T}\|_{F}^{2}\text{.}$ (7)
Now we rewrite it in terms of trace
$\|{\mathbf{L}}\|_{F}^{2}\cdot\|\hat{X}\hat{X}^{T}\|_{F}^{2}=tr({\mathbf{L}}^{2})\text{.}tr((\hat{X}\hat{X}^{T})^{2})\text{.}$ (8)
Because both ${\mathbf{L}}^{2}$ and $(\hat{X}\hat{X}^{T})^{2}$ are positive semi-definite, we have
$tr({\mathbf{L}}^{2})tr((\hat{X}\hat{X}^{T})^{2})\leq tr({\mathbf{L}})^{2}tr(\hat{X}\hat{X}^{T})^{2}\text{.}$ (9)
Finally, combining the above equations, we get the desired result
$nh_{G}\leq\text{tr}({\mathbf{L}})\text{tr}(\hat{X}\hat{X}^{T})\text{.}$ (10)
Proof of Prop. III.3.
Algorithm 1 involves two main steps: computation of the homophily $tr(-{\mathbf{L}}\hat{X}\hat{X}^{T})$, and of the node scores $\mbox{diag}({X}{X}^{T})$. Computation of $tr(-{\mathbf{L}}\hat{X}\hat{X}^{T})$ has complexity $O(dm)$, as it is dominated by computing the matrix-matrix multiplication ${\mathbf{L}}X$ and the subsequent trace computation only requires $O(dn)$. Similarly, the computation of $\mbox{diag}({X}{X}^{T})$ requires $O(dn)$ as we only calculate the diagonal elements of ${X}{X}^{T}$. ∎